Please help. If s varies directly as r^2, and s=64 when r=2. Find s when r=5. Thanks for your help.
Answer by rothauserc(4718)

s varies directly as r^2, then we have
s = kr^2 where k is the proportionality constant
we are given
64 = k(2^2)
k = 64/4 = 16
now we are asked to find s when r is 5
s = 16(5^2)
s = 16 * 25
s = 400